OpenStudy (majorwhipage1):

The equation of line xy is (y-3) = -2/3(x-4) what is the slope of a line perpendicular to xy -2 3 -2/3 3/2

OpenStudy (sshayer):

what is slope of given line?

OpenStudy (majorwhipage1):

Idk thats all the given info

OpenStudy (kgrendel0324):

-2/3

OpenStudy (sshayer):

\[y-y _{1}=m \left( x-x _{1} \right)\]

OpenStudy (sshayer):

slope of line perpendicular to this line is \[-\frac{ 1 }{ m }\]

OpenStudy (sshayer):

now can you tell me the slope of perpendicular line?

OpenStudy (majorwhipage1):

Didn't that guy already say the answer lol

OpenStudy (sshayer):

no,he told only the slope of given line.

OpenStudy (majorwhipage1):

Hmm

OpenStudy (majorwhipage1):

Yeah the topic of slope is insanely confusing

OpenStudy (sshayer):

negative reciprocal. \[=-\frac{ 1 }{ \frac{ -2 }{ 3 } }=1*\frac{ 3 }{ 2 }=\frac{ 3 }{ 2 }\]

OpenStudy (majorwhipage1):

How would you solve after that?